differentiate

f(x) = 2 + |x|2/3

1
Expert's answer
2021-07-20T13:05:48-0400

Solution;

f(x)=2+|x|"^{\frac 2 3}"

Rewrite |x| as "\sqrt{x^2}"

f'(x)="\frac{d(2)}{dx}" +"\frac{d(\sqrt{x^2})^{\frac 23}}{dx}"

Simplify;

f'(x)="\frac{d(2)}{dx}+\frac{d(x^2)^{\frac13}}{dx}"

We have;

"\frac{d(2)}{dx}=0"

Using chain rule;

"\frac{d({x^2})^{\frac13}}{dx}" = "\frac{d(u)^{\frac13}}{du}×\frac{du}{dx}"

Where u=x2

"\frac{d(u)^{\frac13}}{du}" ="\frac13×u^{\frac13-1}" ="\frac{1}{3u^{\frac23}}"

"\frac{du}{dx}"=2x

f'="0+\frac{2x}{3\sqrt[3]{x^4}}"

Since it's cube cube root of "x^4",we don't require an absolute value of x ,since x can be positive or negative. So we simplify as;

f'(x)="\frac{2}{3\sqrt[3]{x}}"